OpenStudy (anonymous):

a bridge is supported by three arches. The function that describes the arches is h(x) = 0.25x^2 + 2.375x where h(x) is the height, in meters, of the arch above the ground at any distance, x, in meters, from one end of the bridge. How far apart are the bases of each arch?

OpenStudy (whpalmer4):

\[h(x) = 0.25x^2+2.375x\]The bases are going to be at the points where \(h(x) = 0\). Can you solve \(0 = 0.25x^2+2.375x\)?

OpenStudy (whpalmer4):

That equation must not be correct. You need a negative term in there somewhere, or it just keeps getting higher....

OpenStudy (whpalmer4):

Are you sure it isn't \(h(x) = -0.25x^2+2.375x\)?

OpenStudy (anonymous):

yes omg you're right so sorry!

OpenStudy (anonymous):

what does that mean!?

OpenStudy (whpalmer4):

Well, it's going to be an inverted parabola. at x = 0, that's the end of the bridge. at the other spot where y = 0, that's the first base.

OpenStudy (whpalmer4):

so, you need to solve the equation to find where h(x) = 0. There are two solutions, although one of them is at x = 0. The other one is the spot where the base goes. Make sense?

OpenStudy (whpalmer4):

This is also the path that a baseball would follow if you threw it up in the air (with no air resistance, that is)

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

whats next

OpenStudy (whpalmer4):

Well, did you solve the equation?

OpenStudy (anonymous):

ahh no, maybe im not getting it-

OpenStudy (anonymous):

where is the second base

OpenStudy (whpalmer4):

Okay, no problem, we'll fix that! :-) So, we want to find the value of \(x\) such that \(h(x) = 0\) to find the other end of our arch. \[h(x) = 0 = -0.25x^2+2.375x\]A little experience tells me that those ugly numbers are all 1/8ths, but even easier is to just multiply through by 1000 and get rid of the decimals altogether. \[0 = -0.25x^2+2.375x\]\[0*1000=-0.25*1000*x^2+1000*2.375x\]\[0=-250x^2+2375x\]We can factor that, or use the quadratic equation, which do you prefer?

OpenStudy (anonymous):

im more familiar with factoring, but whichever is best to teach

OpenStudy (whpalmer4):

Why don't you try factoring that and show me what you get...

OpenStudy (anonymous):

would you mind doing it but showing exactly how you got there, its easier for me to see the proper steps, especially with crazy high numbers : $

OpenStudy (whpalmer4):

Well, do you see any common factors in -250x^2 + 2375x?

OpenStudy (anonymous):

5?

OpenStudy (whpalmer4):

Yes. A whole mess of them, in fact :-) What else?

OpenStudy (whpalmer4):

As a first step, can't you write x(-250x+2375) ?

OpenStudy (anonymous):

25?, 10?

OpenStudy (whpalmer4):

\[-250x^2+2375x = 0\]Factor out \(x\) \[x(-250x + 2375) = 0\]For that to be true, either \(x=0\) or \(-250x + 2375 = 0\) I bet you can solve both of those equations :-)

OpenStudy (anonymous):

ive never factored without a middle number

OpenStudy (whpalmer4):

\(x=0\) is the case for the bank end of the arch. \(-250x + 2375 = 0\) is the case for the other end of the arch.

OpenStudy (whpalmer4):

I'm sure you have, you just don't remember it as such. You maybe haven't solved an equation by factoring where that is true, but I guarantee that you must have factored something like \(a^2 + ab\) when you were first learning how to factor :-)

OpenStudy (anonymous):

uhhh im so confused, what do i do next ?

OpenStudy (whpalmer4):

Solve \[-250x + 2375 = 0\]

OpenStudy (anonymous):

is it 9.5

OpenStudy (anonymous):

i had to use square root factor

OpenStudy (whpalmer4):

No. No square root needed here. \[-250x + 2375 = 0\]Add 250x to both sides \[-250x + 250x + 2375 = 250x\]\[2375=250x\]Now what do you think is the value of \(x\)?

OpenStudy (anonymous):

9.5

OpenStudy (whpalmer4):

Oh, so sorry...9.5 is correct. I wrote it down as 19/2 :-)

OpenStudy (anonymous):

haha thank you soo much!

OpenStudy (whpalmer4):

I'd say I was testing your confidence, but... :-)

OpenStudy (whpalmer4):

Okay, so at x = 9.5, the height of the arch should be 0 again. Let's check: -0.25(9.5)^2 +2.375(9.5) = 0 -0.25*90.25+2.375*9.5=0 -22.5625 + 22.5625 = 0 Success!

OpenStudy (whpalmer4):

Here's a plot of h(x) from x = 0 to x = 10:

OpenStudy (whpalmer4):

So, the bases are 9.5 meters apart, because that's how far it is from x = 0 to x = 9.5.

OpenStudy (whpalmer4):

I'll quickly do it with the quadratic formula, just for illustration: we had \[h(x) = -0.25x^2+2.375x\]so our quadratic would be\[-0.25x^2+2.375x = 0\]giving us \(a=-0.25, b = 2.375, c = 0\). Our solutions would be thus\[x = \frac{-2.375\pm\sqrt{2.375^2-4(0)(-0.25)}}{2*(-0.25)} = \frac{-2.375\pm2.375}{-0.5} = 0, 9.5\]